A rectangular land of 90 hectares.
If the width if 2/5 of the length.
What is the perimeter?

Answer by Cromlix(4381)

Hi there,
Make length = x
the width is therefore 2/5x
Area = length x width
90 = x(2/5x)
Reverse equation
2/5x^2 = 90
2x^2 = 90 * 5 [* means times]
2x^2 = 450
x^2 = 450/2
x^2 = 225
x = 15
Length = 15 hectares
Width = 6 hectares
Perimeter = 2 * length + 2 * width
Perimeter = 2 * 15 + 2 * 6
Perimeter = 42
If you want your measurements in metres
90 hectares = 900,000 m^2
Area = length x width
900000 = x(2/5x)
Reverse equation
2/5x^2 = 900000
2x^2 = 900000 * 5 [* means times]
2x^2 = 4500,000
x^2 = 4500000/2
x^2 = 2250000
x = 1500
Length = 1500 hectares
Width = 600 hectares
Perimeter = 2 * length + 2 * width
Perimeter = 2 * 1500 + 2 * 600
Perimeter = 4200 metres
Hope this helps :-)